OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3).

OpenStudy (anonymous):

slope of the required line = -1/2 = m now use y-y1 = m(x-x1) where m = -1/2, x1= 6 and y1= 3